OpenStudy (anonymous):

If gas costs $3.65 per gallon at a local service station, how many cubic feet of gas can a customer buy with $40 ?

OpenStudy (precal):

40/3.65=10.95 gallons

OpenStudy (anonymous):

how far can the customer drive if the car gets 4 miles per cubic foot of gas?

OpenStudy (precal):

4 miles per gallon??? Horrible mileage.....

OpenStudy (anonymous):

the first answer was rejected, the 10.95

OpenStudy (anonymous):

they want the answer in ft^3, i guess

OpenStudy (precal):

weird.... I was just thinking of it like in the real world. I have $40 dollars and I can only by approx 10.9 gallons at 3.65 maybe you need to input it as 10.9 (do you have instructions on what place to round to?)

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

well can you help me with 4 other questions please

OpenStudy (anonymous):

i have this The compound interest formula is F = P(1 + i)^n where F is the total amount due, P is the initial cost or amount of money, i is the interest rate, and n is the number of periods interest is computed. A student purchases a $1300 laptop computer for school. The deal does not require payment until 2 years later. If interest is computed monthly, how much does the student have to pay at the end of 2 years? The monthly interest rate is 1.8 percent.

OpenStudy (precal):

ok I will try, what do you have?

OpenStudy (precal):

P=1300 i=.018 n = 12 because it is monthly but n = 24 because of 2 years F=1300(1+.018)^24 compute F

OpenStudy (anonymous):

did you get 1994?

OpenStudy (precal):

and 76 cents

OpenStudy (precal):

That would be a bad deal, that means that person is willing to pay 694.76 to borrow 1300

OpenStudy (precal):

yes 1994.76

OpenStudy (anonymous):

my damn software rejected the answer

OpenStudy (precal):

I still think the directions somewhere has to say how to input the solutions. Decimals are tricky. I am given you real world solutions but if the instructions are to round to the nearest whole number 1995 would be the only accepted answer.

OpenStudy (anonymous):

ok what about this one A copper pipe, used for plumbing from the water main to a house, is 5 meters long and has a circular cross-section 3.2 cm in diameter. There is a 2.4 cm diameter hole in the center of the cross-section for the entire length of the rod. What is the volume of copper in the rod in cubic inches?

OpenStudy (precal):

were you given a picture? Upload it by attaching the file.

OpenStudy (anonymous):

yea theres the picture for the last problem

OpenStudy (precal):

V=Bh B is a circle so B is pi(r^2) \[V=\pi r^2h\]

OpenStudy (anonymous):

so v=pi(5^2)(3.2)?

OpenStudy (precal):

picture stated 1 gallon = 231 cubic inch we sort of needed that info earlier

OpenStudy (anonymous):

oh sorry

OpenStudy (precal):



OpenStudy (precal):

original solution is 10.95*231=2529.45 cubic inches

OpenStudy (precal):

 I think you will have a circle within a circle, but you should not be using diameter. Take half of the diameter and that is your radius

OpenStudy (anonymous):

3.2/2=1.6, is that right

OpenStudy (precal):

yes if d=3.2 then r = 1.6

OpenStudy (anonymous):

so is h the 2.4?

OpenStudy (precal):

no 5 is the height (it you stand the pipe up you will see that it is 5)